### v2

# SonicVisionLM: Playing Sound with Vision Language Models

Abstract
There has been a growing interest in the task of generating sound for silent videos, primarily because of its practicality in streamlining video post-production. However, existing methods for video-sound generation attempt to directly create sound from visual representations, which can be challenging due to the difficulty of aligning visual representations with audio representations. In this paper, we present SonicVisionLM, a novel framework aimed at generating a wide range of sound effects by leveraging vision language models. Instead of generating audio directly from video, we use the capabilities of powerful vision language models (VLMs). When provided with a silent video, our approach first identifies events within the video using a VLM to suggest possible sounds that match the video content. This shift in approach transforms the challenging task of aligning image and audio into more well-studied sub-problems of aligning image-to-text and text-to-audio through the popular diffusion models. To improve the quality of audio recommendations with LLMs, we have collected an extensive dataset that maps text descriptions to specific sound effects and developed temporally controlled audio adapters. Our approach surpasses current state-of-the-art methods for converting video to audio, resulting in enhanced synchronization with the visuals and improved alignment between audio and video components. Project page: https://yusiissy.github.io/SonicVisionLM.github.io/
Figure 1: A model implements the automatic detection of in-picture voice generation and accepts the user’s editing of this paper and time in the off-screen section.On-screen sound means sound originate from visual information.Off-screen sound not directly happened on the screen.
Figure 2: SonicVisionLM’s framework. SonicVisionLM presents a composite framework designed for the automatic recognition of on-screen sounds coupled with a user-interactive module for off-screen sounds editing. The blue dashed box and arrows in the figure represent the visual automation workflow: First, a silent video goes into the system where it figures out what’s happening (text)and when(time). Then, this info is used to make the sounds that match what’s on screen. The purple dotted box and arrows show how users can add their own changes to create off-screen sounds.
1 Introduction
The sound effects artists work with different types of sounds, some you could see on-screen (like footsteps or a car passing) you couldn’t see but hear (like background noises that make the video feel more real and add to the story). On-screen sounds match what’s happening in the video, while off-screen sounds help set the mood and give extra information. Making sounds for videos is a key part of video production, however, it could take lots of labor for the artists. As a result, the video-sound generation task has gained significant attention.
Although the recent approaches have made great efforts, the video-sound generation task is still challenging. For on-screen sounds, achieving semantic relevance and maintaining temporal synchronization continues to be a complex issue. For off-screen sounds, it is hard to edit. Current methods [31, 13, 24] primarily focus on the visual content to generate the corresponding sound, a subset of these methods [8, 7] takes into account the aspect of editability. However, the alignment between video features and audio features is quite difficult, leading to deficiencies including 1) incorrect sound meanings, and mismatched timing, 2) monotonous sound effects, and lack complex scenarios. Both lead to unsatisfactory results in video-sound generation task.
To solve the above deficiency, we propose a novel framework, named SonicVisionLM, as shown in Fig.1. SonicVisionLM is proposed by introducing three key components, including video-to-text, text-based interaction, and text-to-audio generation. First, the video-to-text component focuses on generating sound effects for on-screen events. In this step, VLM is employed to identify pertinent sound descriptions from the input silent video. Following this, a temporal detection network is trained to extract specific text-based interact temporal information from the video. A key innovation within this framework is the design of a time-conditioned embedding, which is utilized to guide an audio adapter. After that, the text-based interaction component allow users to change the text and timestamps from a previous video-to-text component, or to input new corresponding text-timestamp pairs for personalized sound design. Finally, the text-to-audio generation component accept the text and timestamp conditions, and then input them into the LDM and adapter to generate diverse, time-synchronized, and controllable sounds. Simultaneously, We collect a text-to-single-sound dataset, named CondPromptBank, for sound effects caption and timing cues, comprising over ten thousand data points, covering 23 different categories, CondPromptBank can generate high-quality and divers sound. The main contributions of this work are:
  • We propose a novel framework called SonicVisionLM and collect a dataset CondPromptBank specifically for training a time-controllable adapter. This ensures the generated sound aligns perfectly with our text input and maintains precise timing control.
  • We introduce three pioneering components: video-to-text, text-based interaction, and text-to-audio generation. This unique combination facilitates the automatic recognition of on-screen sounds while enabling user customization of off-screen sounds.
  • The proposed framework achieves state-of-the-art result on both conditional and unconditional video-sound generation tasks.The conditional task can be noticeably enhanced in all metrics. (IoU: 22.4$\rightarrow$39.7)
2 Related Work
We roughly divide audio generation tasks into two categories. The domain of Text-to-Audio generation encompasses various subfields, including Text-to-Speech (TTS) and Text-to-Music (TTM). State-of-the-art TTS models like FastSpeech2 [28] and NaturalSpeech [32] have made significant progress, producing speech quality nearly indistinguishable from human speech. In the TTM domain, MusicLM [1], Noise2Music [11], MusicGen [6], and MeLoDy [20] are aimed to generate music segments from textual descriptions, bringing significant innovation to the field of music composition and synthesis. Models like AudioGen [18], AudioLDM [23], Tango[9], and Make-an-Audio [12] focus on universal audio generation modeling. AudioGen [18]treats audio generation as a conditional language modeling task, while the other three models employ latent diffusion methods to accomplish sound generation. However, due to the comprehensive large-scale datasets, existing models lack the direct capability to generate sound effects for videos. Therefore, we created a text-to-single-sound dataset to help the model produce high-standard sound effects. In Video-to-Sound Generation task, SpecVQGAN [13]employs a Transformer-based autoregressive model, utilizing extracted ResNet50 or RGB+Flow features for sound generation. Im2Wav [31]uses a dual-transformer model conditioned on CLIP features for sound generation. CondFoleyGen [8]and VARIETYSOUND [7]have introduced tasks for controllable timbre generation. Diff-foley[24] uses contrastive audio-visual pretraining to align audio and visual features However, the results generated by these methods have poor audio-visual synchronization, obvious noises, and uneditable. Unlike the previously mentioned works, our model not only audio-visual synchronization and diversity, but also allows for personalized edits from users. Our model provides a more comprehensive sound solution for video production.
Diffusion Model. have been explored for both mel-spectrogram generation [27, 4], and waveform generation [19, 21, 3]. However, the iterative generation process can be slow for high-dimensional data. AudioLDM[23], Make-An-Audio[12], and Tango[9], which trained in a continuous latent space. However, the control of LDM in audio generation tasks still falls short of achieving satisfactory results. In this paper, we attempt to control the timing of sound generation and ensure audio-visual synchronization.
Vision Language Models. Recently, GPT-4 has showcased advanced multi-modal abilities and inspired vision-language LLMs. Vision-LLM [35] and LLaVA[33] are aligning image inputs with large language models Vicuna[5] exhibit similar multi-modal capabilities. Recent developments in this field include MiniGPT-v2[2]. Kosmos-2 [26] demonstrating the ability of multi-model LLMs to perform visual grounding. In this paper, we first introduce the VLM to the audio generation task.
3 Method
3.1 Overview
In this section, we introduce the framework of SonicVisionLM, as shown in Fig. 2. Before delving into the specific design details, we first provide a brief overview of the preliminary knowledge (Section 3.1). Then, we introduce the Visual-to-Audio Event Detection Module (Section 3.2), which obtains textual descriptions of on-screen sounds through VLMs. Subsequently, we present the Sound Event Timestamp Detection Module (Section 3.3), designed to accurately detect the timing cues through network architecture. Finally, we introduce the proposed time-controllable adapter as an extension of the audio diffusion model (Section 3.4), enabling the generation of multiple sounds that are semantically coherent and temporally aligned.
3.2 Preliminaries
Audio Diffusion Model. The text-prompt encoder encodes the input description $\tau\in R^{L\times d_{t}}$ of the sound, where $L$ is the token count and $d_{t}$ and is token-embedding size.
The latent diffusion model (LDM) to construct the audio prior $z_{0}$ with the guidance of text encoding $\tau$. This essentially reduces to approximating the true prior $q\left(z_{0}\mid\tau\right)$ with parameterized $p\left(z_{0}\mid\tau\right)$. LDM can achieve the above through forward and reverse diffusion processes. The forward diffusion is a Markov chain of Gaussian distributions with scheduled noise parameters $10<\beta_{1}<\beta_{2}<\textperiodcentered\textperiodcentered\textperiodcentered<\beta_{N}<$ 1 to sample noisier versions of $z_{0}$, where $N$ is the number of forward diffusion steps, $\alpha_{n}=1-\beta_{n}$, and $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{n}$ :
$q\left(z_{n}\mid z_{n-1}\right)$ $=\mathcal{N}\left(\sqrt{1-\beta_{n}}z_{n-1},\beta_{n}\mathbf{I}\right),$ (1)
$q\left(z_{n}\mid z_{0}\right)$ $=\mathcal{N}\left(\sqrt{\bar{\alpha}_{n}}z_{0},\left(1-\bar{\alpha}_{n}\right)\mathbf{I}\right),$ (2)
where the noise term $\epsilon\sim\mathcal{N}(0,I)$. The final step of the forward process yields $z_{N}\sim\mathcal{N}(0,I)$.
The reverse process denoises and reconstructs $z_{0}$ through text-guided noise estimation $(\hat{\epsilon}_{\theta})$ using loss:
$\mathcal{L}_{DM}=\sum_{n=1}^{N}\gamma_{n}\mathbb{E}_{\epsilon_{n}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),z_{0}}\left\|\epsilon_{n}-\hat{\epsilon}_{\theta}^{(n)}\left(z_{n},\tau\right)\right\|_{2}^{2},$ (3)
After LDM is trained, we generate audio latent by sampling through the reverse process with $z_{N}\sim\mathcal{N}(0,I)$, conditioned on the given textual representation $\tau$, with the following reverse dynamics:
$p_{\theta}\left(z_{n-1}\mid z_{n},\tau\right)$ $=\mathcal{N}\left(\mu_{\theta}^{(n)}\left(z_{n},\tau\right),\tilde{\beta}^{(n)}\right),$ (4)
$\mu_{\theta}^{(n)}\left(z_{n},\tau\right)$ $=\frac{1}{\sqrt{\alpha_{n}}}\left[z_{n}-\frac{1-\alpha_{n}}{\sqrt{1-\bar{\alpha}_{n}}}\hat{\epsilon}_{\theta}^{(n)}\left(z_{n},\tau\right)\right],$ (5)
$\tilde{\beta}^{(n)}$ $=\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha}_{n}}\beta_{n}.$ (6)
The noise estimation $(\hat{\epsilon}_{\theta})$ is parameterized with U-Net framework [29] with a cross-attention component to include the text guidance $\tau$.
Thereafter the decoder of audio VAE [17] constructs a mel-spectrogram feature from the latent audio representation $\hat{z}_{0}$. This mel-spectrogram feature is convey to a vocoder to generate the final audio.
3.3 Visual-to-Audio Event Understanding Module
The diversity of sound is influenced by various factors, including the source of the sound, actions, the environment, and more. Common methods of naming sound effects are also based on these factors. Inspired by the widespread use of visual language models[22, 36, 38], we chose miniGPT-v2[2]to process visual information and generate descriptions of sounds. Due to miniGPT-v2 is originally designed for single-image understanding, recognizing the limitation of single-image in conveying dynamic information, We follow the LLaMA-2[33] conversation template design and adapt it for the multi-modal instructional template. we made adjustments to accept the four random video frames feature and incorporated temporal cues in the instructions. The adapted template is defined as follows:
$First,\ \ \ \ \ \ \ \ <Img><ImageFeature><Img>.\\ \ Then,\ \ \ \ \ \ \ \ <Img><ImageFeature><Img>.\\ \ Afterthat,<Img><ImageFeature><Img>.\\ \ Finally,\ \ \ \ \ <Img><ImageFeature><Img>.\\ \left[Task\ \ Identifier\right]\ \ \ Instruction$
3.4 Sound Event Timestamp Detection Module
We use the sound event timestamp detection module is used to detect the timestamp of the sound event in the video inspired by the CondFoleyGen[8]. Instead of using a hand-crafted approach to transfer sounds from the conditional example, We train a ResNet (2+1)-D18[34] visual network to capture the timestamp of on-screen sound events. We pre-trained this network using a dataset comprising videos and sound timing data. The dataset generates the ground truth timestamp with the timing label that is detected by aligning it to the closest frame. We use a binary cross-entropy loss, penalizing time predictions that occur at incorrect times. Since each input video can have more than one sound event, we set the weight of each onset in the loss from the video equally according to the total number of onsets in the input so that the weight sums to one. Subsequently, we input the silent video frame sequence $V_{f}$ into the detection network, and the output is a one-hot vector $V_{ct}$ of predictions using a fully connected layer after pooling (one prediction per input frame). This vector serves as input for the subsequent Time-controllable Latent Diffusion Model. The whole process and network structure are shown in Fig.3.
Figure 3: Sound Event Timestamp Detection Module.The network organizes the video’s features to create a special heat vector. This vector is as long as the number of frames in the video. The white parts in it show where there’s sound, and the black parts show where there’s no sound.
3.5 Time-controllable Latent Diffusion Model
In our experimentation, we observed that the results generated were semantically inaccurate and temporally unsynchronized. It would happen when such methods employed audiovisual datasets for training end-to-end models. This problem hinders the widespread application of this task in reality. We posit that to solve effectively aligning visual features with corresponding audio features could solve this problem. Therefore, we propose using timestamps to control the text-to-sound generation model, which improves sound quality, diversity, and synchronization.
In the visual domain, many works are based on the ControlNet architecture, which can finer control the generation of images or videos by manipulating the input image conditions of neural network blocks. However, unlike the visual domain’s inherent intuitiveness, audio features exhibit temporal continuity and are inherently more complex and abstract. Consequently, the selection of appropriate audio features to guide the generation process presents a significant challenge. In this paper, we propose a embedding called audio time-condition embedding $A_{ct}$, which involves the following steps: Firstly, we use function $F(\cdot)$ extract the start $x_{start}$ and end points $x_{end}$ of sound segments from the audio $X_{a}$ and expand them into one-hot vectors $V_{ct}$of the same length as the audio, where these one-hot vectors indicate the presence of sound in each audio frame.
$x_{start},x_{end}=F\left(a_{c}\right)$ (7)
$V_{ct}=\begin{cases}1\ \ \ ,if\ \ t\ \ in\ \ \ [x_{start},x_{end}]\\ 0\ \ \ ,else\end{cases}$ (8)
Subsequently, based on these one-hot vectors $V_{c}t$, we fill all channels of the mel-spectrogram with maximum values at the corresponding time frames Finally, we obtain mel feature embeddings $A_{ct}$ through a VAE encoder $E_{a}$.
$a_{ct}=E_{a}\left(\operatorname{max}\left(\operatorname{mel}\left(a_{c}\right),V_{ct}\right)\right)$ (9)
we have taken inspiration from ControlNet[39], designed a network architecture,called Time-controllable Adapter. Then, the audio time-condition embedding $A_{ct}$ is with the text embedding $\tau$and target audio embedding, fed into the neural network block to train the time-controllable adapter together. For brevity, we denote the encoder as $F$ , the middle block as $M$, and the decoder as $G$, with $f_{i}$ and $g_{i}$denoting the output of the $i-th$ block in the encoder and decoder, and $m$ denoting the output of the middle block, respectively. Specifically, we fix the weights of Tango and copy the structures and weights of the encoder and middle block, designated as $F^{\prime}$ and $M^{\prime}$respectively. Thereafter, we incorporate the information from the time-controllable adapter during the decoding process. The output of the time-controllable adapter is added to the output of the corresponding block of the decoder. To achieve it, we ensure that all other elements are kept frozen while modifying the input of the $i-th$ block of the decoder as:
[…]
where $zero$ represents one zero convolutional layer whose weights increase from zero to gradually integrate control information into the main diffusion model. This approach was adopted to preserve the audio generation capabilities of the large-scale model, which was trained on billions of audio-text pairs. Simultaneously, it enabled the model to comprehend the guidance provided by the time control embedding, resulting in temporally controllable outcomes.
[…]
$*{}^{*}$[8] 15.5 21.6 67.0 35.5 23.3
CondFoleyGen(codebook) $*{}^{*}$[8] 16.3 19.2 68.6 37.4 22.4
Ours-small 21.5 19.4 77.03 27.8 35.6
Ours-full 46.7 27.6 78.1 43.8 39.7
Table 1: Conditional Generation Task Qualitative Results. CLAP-top1 is a metric that evaluates the model’s conditional controllability. It calculates the percentage of times the sound samples generated by the four models rank in the top 1 according to CLAP rankings, divided by the total number of samples. Onset Acc and Time Acc are both metrics based on the number of sound occurrences, with Onset Acc focusing on the onset count, and Time Acc on the count of time intervals. We measure the average accuracy of predicting onsets within 0.1 seconds of ground truth to assess the timing of generated onsets. IoU is calculated by computing the intersection and union of these vectors. These metrics collectively allow us to comprehensively evaluate the accuracy of the generated sounds in terms of both timing and content. $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance. The five metrics all measured in percentage.
4 Experiments
4.1 Experiment Settings.
Figure 4: Conditional Generation Task Qualitative Results. The red dashed boxes are the conditional audio inputs and generated results for the CondFoleyGen task, and the blue dashed boxes are the conditional text inputs and generated results corresponding to SonicVisionLM.
Dataset. Considering that Tango is a large model that incorporates various types of audio data, including speech, sound effects, and music, we have developed a high-quality single sound effects dataset called CondPromptBank. This dataset is specifically designed for training a time-controllable adapter, which allows us to align the generated sound with our task requirements while ensuring precise timing control.
CondPromptBank consists of 10,276 individual data entries. Each entry includes a high-quality sound effect, its corresponding sound effect title, and the starting and ending timestamps for when the sound is emitted. The duration of these sound effects is typically 10 seconds or shorter. All the data in CondPromptBank is sourced from freely available sound effect libraries and websites. During collection, we focused on 23 common categories of sound effects and manually filtered out low-quality data with high levels of noise and mixed sources.
To provide comprehensive textual descriptions, we further annotated the sound effect text labels with fine-grained information based on sound characteristics. Each sound effect text label now includes details about the single audio source and the associated action. We believe that this division into single audio sources is crucial because videos, as compared to text, provide a more concrete expression of content, demanding a greater accuracy in sound representation.
Implementation details. To train our complete model, we first train the adapter on CondPromtBank, then train the timestamps detection net using a large video-audio-time dataset. We trained the adapter for approximately 200 epochs with a batch size of 32 and a learning rate of $3.0\times 10^{-5}$ using Adam [16]. Our modal has two versions: small and full, the difference between them is the versions of pre-trained parameters used for Tango. We trained the timestamps detection net for 70 epochs with a batch size of 24 and a learning rate of $1.0\times 10^{-5}$. The training of the adapter takes approximately 5 days and the training of the the timestamps detection network requires approximately 1 day using 1 NVIDIA A6000 GPU.
Metric
Dataset Greatest Hits Countix-AV
[…]
$*{}^{*}$[13] 6.80 82.4 2.17 25.8 7.39 34.1 5.29 34.9
$*{}^{*}$[24] 6.43 31.3 4.75 24.7 8.09 17.9 15.3 32.6
Ours-small 6.46 31.9 3.88 36.6 10.0 21.7 15.1 37.5
Ours-full 4.67 24.9 3.26 39.5 9.71 19.7 12.7 42
Table 2: Unconditional Sound Generation Quantitative Results. IS assesses the quality and diversity of generated samples, FID measures distribution-level similarity, and MKL measures paired sample-level similarity. $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance. IoU metric is measured in percentage.
4.2 Conditional Generation Task Results
We tested our model using the Greatest Hits[25] dataset, which has 977 videos of drumsticks interacting with different objects, lasting 11 hours. This dataset is divided into two types of actions and 17 types of materials. This detailed categorization helps check if our model can change sound types based on these details but still match the target action. We used the same test settings as CondFoleyGen[8] and compared our results with theirs.
Figure 5: Unconditional Generation Task Qualitative Results. The left example is from CountixAV, and the right one is from Greatest Hits. We’re comparing them side by side. The dashed box highlights examples of both good and bad results we generated.
Evaluation Metrics. For the conditional generation task, we use the following five objective metrics to evaluate the performance of the model: CLAP-top1, Onset Acc[8], Onset AP[8], Time Acc, and IoU.
Qualitative Results. As seen in Tab.1, SonicVisionLM-full demonstrates superior performance over CondFoleyGen across all metrics. It got a CLAP-top1 score of 46.7%, 30% higher than any versions of the CondFoleyGen model This result suggests that it’s better at matching sounds to text prompts. In audio-visual sync, our model outperforms CondFoleyGen significantly, achieving an Onset AP score that is 11% higher and an IoU score that is 16% higher. These results underscore our model’s enhanced ability to generate sounds that are more accurately synchronized with the input video in terms of timing. Our model outperforms CondFoleyGen in onset and timing accuracy, showing improvements of 6% and 8.3%, respectively. This underscores its precision in detecting sound events and its effectiveness in aligning visual and audio elements.
We think CondFoleyGen relies on the audio-visual synchronization module[14] to improve time accuracy, which needs a large number of samples to re-ranking the sounds. In contrast, our model gets higher synchronization during the generation process, with fewer samples.
Quantitative Results. As shown in Fig.4 (row 7,8,9), we compare the timestamped positional distance to the target sound. Our results closely match the target sound, demonstrating a high degree of accuracy. However, CondFoleyGen often produces results with the wrong number of sounds and with a large difference in the positional distance This indicates that our model has a higher visual-audio synchronization than CondFoleyGen. As shown in Fig.4 (column 3,7), we compare the sound shape to the conditional sound. Even though our model is not trained on the Greatest Hits dataset, our results are still very similar to the conditional sound. As shown in Fig.4 (column 1,7), The mel-spectrograms of the CondFoleyGen results do not produce sound that are similar to, and sometimes blurred.
4.3 Unconditional Generation
To evaluate the task of unconditional sound generation, considering that our model has not been trained on audio-visual datasets, we have chosen to perform quantitative evaluation and qualitative evaluation on two datasets: Greatest Hits[25] and CountixAV[40], which are zero-shot tasks for all models. We use two state-of-the-art condition V2A models: SpecVQGAN [13] and DIFF-FOLEY[24] as baselines.
Evaluation Metrics. For objective evaluation, we have employed three metrics as [13, 24]: Inception Score (IS)[30], Frechet Distance (FID)[10], and Mean KL Divergence (MKL)[13].
For subjective evaluation, As the [8], we conduct user evaluations for the three critical components: overall audio quality (OVL), alignment with the input video (REL), and time synchronization (Time-sync).
[…]
$*{}^{*}$[13] 37 25 31
$*{}^{*}$[24] 48 64 58
Ours 75 69 87
Table 3: Subjective Results. Following [8], we invited the 300 evaluators to judge 30 randomly selected audio samples from three perspectives: OVL, REL, and Time-sync. Ratings were scored on a scale from 1 to 100. All evaluators were proficient in English. The scoring scale is 1-100 and results are averaged out. $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance.
Quantitative Results. As shown in Tab.2, our model notably surpasses DIFF-FOLEY and SpecVQGAN on the Greatest Hits dataset in terms of MKL and FID metrics, with respective improvements of 1.76, 6.4, and 2.13, 57.5. This signifies our model’s superior ability to generate sounds that align closely with the ground truth. On the CountixAV dataset, SonicVisionLM-small is only marginally behind the best result in the IS metric by 0.2, suggesting a commendable sound quality produced by our model. Despite our model’s slightly lower performance in the MKL metric compared to the baseline, we view this not as a drawback. Unlike the Greatest Hits dataset, which was recorded in high quality with specialized recording equipment, the CountixAV dataset is sourced from diverse YouTube videos, and often includes sounds marred by low-quality background noise or non-homogenous sound sources. The MKL and FID metrics, focusing on similarity to ground truth, which could explain why SpecVQGAN, despite primarily generating noisy sounds, excels in the MKL metric. The IoU metrics show that our model makes sound at the right times and keeps quiet at other times. To further argue our point, we conducted a subjective assessment of three aspects of OVL, REL, and Time-sync. We found that our model performed significantly better than the baseline models in all metrics, as shown in Table.3
Qualitative Results. To compare the time points more visually, we denoised the CountixAV dataset. As we can see from Fig.5, our generated sounds are more clear than the baseline models. On top of that, our model can capture finer-grained visual transformations, capturing the signal despite sudden changes in the frequency of movement and producing the correct sound. in contrast, SpecVQGAN results have only noise throughout the sound. DIFF-FOLEY sometimes can generate sounds without noise, but it is not aligned with the visual-sound synchronization.
[…]
$*{}^{*}$[9] 44.9 11.5 3.47 41.3
Ours-small 26.2 2.05 2.30 71.6
$*{}^{*}$[9] 44.3 10.87 3.14 39.3
Ours-full 25.4 3.19 2.32 65.7
Table 4: Time Condition Ablation $*{}^{*}$ ”denotes the data sourced from the official code and is based on experiments conducted with our local configurations. underline denotes the worst performance. boldface denotes the best performance. IoU metric is measured in percentage.
Figure 6: Conditioning-scale Ablation Study. In the chart, the higher the indicators are, the better they are. The brightness of the color also represents a good or bad performance.
4.4 Ablation Study
Time condition. we compared our model with the tango[9]. Evaluation Metrics. In the ablation study, We employed objective metrics from Tango[9], FAD [15] and Distance (FD)[23] measure the distribution difference between generated audio and real audio without the need for any reference audio samples. KL divergence [37, 18] assesses the similarity between them based on the broad concepts present in the original and the generated audio signals.
Result. As shown in Tab.4, both in the small and full versions, our model outperforms tango in terms of IoU, which validates the effectiveness of the time-controllable adapter.
Conditioning-scale. As seen in the Fig.6. In the range from 0.6 to 2.0 on the conditioning scale, the sound quality improves as it lines up better with the IoU. But at 3.0, control goes down a lot. The best mix of good sound quality and steadiness is found at 2.0.
4.5 Multi-soundtracks generation
As shown in Fig.7, SonicVisionLM works like a sound designer in traditional video post-production. First, it carefully looks at the visual information to find the key sounds needed. For example, it picks out the hissing of cooking meat and the clinking of pots. Then, if the story needs it, the sound designer adds off-screen sounds. This might include extra kitchen sounds like distant chopping, simmering pots, or the buzz of kitchen machines, making it feel more real. Lastly, SonicVisionLM puts these audio tracks in the right place on the video timeline. This makes sure the sounds match up perfectly with what’s happening in the video, creating a smooth audio-visual experience. More detailed results can be found in the Supplementary Materials.
Figure 7: Multi-soundtracks generation example. In the video, the blue and green masks show different soundtracks. You can notice that the timing of on-screen sound in the video (like actions happening) matches when they happen in the frame. But the sounds that come from off-screen (not seen in the video) don’t have precise timing.
5 Conclusion
In this paper, we propose SonicVisionLM, which utilizes the capabilities of powerful vision language models (VLMs). When provided with a silent video, SonicVisionLM first identifies events within the video using a VLM to suggest possible sounds that match the video content. SonicVisionLM demonstrates outstanding performance in both conditional and unconditional generation tasks. Additionally, we tested its efficacy in the post-production process, focusing on automatic recognition of on-screen sounds and personalized editing for off-screen sounds. Extensive experiments show superior performance of our method.
Limitations. While SonicVisionLM has achieved satisfactory results, refinements in the visual understanding and timestamp detection parts are still required. When it comes to various multimedia contexts, expanding the diversity and range of audio generation control are still the key demands of applications. Thus, the model should enrich the audio-visual experience and broaden applicability in future attempts.
Supplementary Material
6 Timestamp Detection Module Precision experiment
Our timestamp detection model shows promising performance, with an Average Precision(AP) of 0.80 and an accuracy of 0.72 on the Greatest Hits test set, and even higher results on the validation set with an AP of 0.92 and accuracy of 0.82. However, its performance on the CountixAV test set is comparatively lower, achieving an AP of 0.52 and accuracy of 0.52. This discrepancy likely stems from the complexity of sound sources in our training dataset, leading to potential inaccuracies in ground truth (GT). Such labeling challenges can adversely affect recognition accuracy, particularly in scenarios involving non-static footage.
7 Subjective Results Details
We provide the screenshot of the main evaluation page that the participant will see during the test in Fig.8.
Figure 8: The evaluation page for the unconditional generation task. we display a screenshot of the primary test interface that participants will encounter. Each participant is required to input three scores for the presented content. Upon clicking the ’Next’ button, they will be directed to the subsequent video in the test sequence.
Figure 9: Screenshot of the conditional generation task section in the demo video.
8 Additional Results
Figure 10: Screenshot of the unconditional generation task section in the demo video.
Figure 11: The process of video-to-text and text-based interaction components in multi-track generation tasks.
Conditional Generation Results As shown in Fig.9, the left column represents the target video, the middle column showcases the control conditions for CondFoleyGen and our model, while the right column displays the generated results. We provide 6 examples for the previously mentioned conditional generation task, and the corresponding audio outcomes can be observed in the demo video.
Unconditional Generation Results As shown in Fig.10, we compare a comparison of the results generated by GT, SpecVQGAN, DIFF-FOLEY, and our model. We provide 6 examples from CountixAV, and the audio samples are available in the demo video.
Multi-soundtracks Generation Results As shown in Fig.11, we segment the video into different shots firstly. Shots with clear actions are directly processed by the VLMs to obtain corresponding sound effect descriptions and their spatial positioning within the video. These descriptions are used to generate on-screen sound. Other shots accept user editing and are then fed into the LDM to produce off-screen sounds.We provide a simple and a complex case respectively in the demo video.